If a company wishes to raise R2 million through a rights issue at a subscription price of R100 each, while its 20 000 outstanding shares trade at R 150 apiece, what would the theoretical value of its shares after issue be?
1
Expert's answer
2020-08-17T18:10:10-0400

Before issue ;the company had 20,000 outstanding shares trading at R150 per share.


The company wants to raise R 2 million through a rights issue at a subscription price of R 100 each hence the number of rights is "\frac{2,000,000} {100}" =20,000 shares.


Thus the theoretical value(after issue) =

"\frac{Market value of shares prior to rights issue +Cash raised from rights issue} 
{Number of shares after rights issue}"


"\frac{(20,000 ×150) +2,000,000} {20,000+20,000}" ="\frac{5,000,000}{40,000}" =125


=R125